RhondaSommer (rhondasommer):

A compound contains 67.1 percent zinc and the rest is oxygen. What is the empirical formula of the compound? ZnO Zn2O ZnO2 Zn2O4

RhondaSommer (rhondasommer):

ZnO right?

OpenStudy (sweetburger):

No Zn is only 65% mass of that compound. Zn2O4 has Zn to be 67.1% by mass

RhondaSommer (rhondasommer):

how does the "2" change the mass?

OpenStudy (sweetburger):

65+65+16+16+16+16= 194. so the mass of Zn/ mass of Zn2O4 = 130/194 =67%

RhondaSommer (rhondasommer):

woah wait a sec. were did all those numbers come from?

OpenStudy (sweetburger):

1 Zn + 1 Zn + 1 Oxygen + 1Oxygen + 1 Oxygen + 1 Oxygen

RhondaSommer (rhondasommer):

Oh!

RhondaSommer (rhondasommer):

I C I C